OpenStudy (anonymous):

A line passes through the point and has a slope of . What is the equation of the line?

OpenStudy (anonymous):

point (0,-3/2) slope 1/2

OpenStudy (abb0t):

slope = m = \(\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (abb0t):

to find point, use: \((y_1-y)=m(x_1-x)\) where y and x are your given points.

OpenStudy (anonymous):

not what i asked but

OpenStudy (abb0t):

to find equation*

OpenStudy (anonymous):

Whats the answer dude i need to know so i dont fail a quiz

OpenStudy (abb0t):

Just plug it in and you'll get your answer, "dude".

OpenStudy (anonymous):

LOL what a joke of a human being.

OpenStudy (abb0t):

watch the 'tude, mortal.

OpenStudy (anonymous):

Are you dumb?